OpenStudy (anonymous):

Expand (tan theta + sec theta)^2

OpenStudy (anonymous):

\[ \left( \tan \theta + \sec \theta \right)^2\]

OpenStudy (anonymous):

Step-by-step, please check me, but I think this is good: \[(\sec(\theta)+\tan(\theta))^2\] \[(\sec(\theta)+\tan(\theta)) (\sec(\theta)+\tan(\theta))\] \[\sec(\theta) (\sec(\theta)+\tan(\theta))+\tan(\theta) (\sec(\theta)+\tan(\theta))\] \[\sec(\theta) (\sec(\theta)+\tan(\theta))+(\sec(\theta)+\tan(\theta)) \tan(\theta)\] \[\sec(\theta) \sec(\theta)+\sec(\theta) \tan(\theta)+(\sec(\theta)+\tan(\theta)) \tan(\theta)\] \[\sec(\theta) \sec(\theta)+\sec(\theta) \tan(\theta)+\sec(\theta) \tan(\theta)+\tan(\theta) \tan(\theta)\] \[\sec(\theta) \sec(\theta)+2 \sec(\theta) \tan(\theta)+\tan(\theta) \tan(\theta)\] \[\sec^{1+1}(\theta)+2 \sec(\theta) \tan(\theta)+\tan^{1+1}(\theta)\] \[\sec^2(\theta)+2 \sec(\theta) \tan(\theta)+\tan^2(\theta)\] If you have questions as to what I did on a particular step please ask away!

OpenStudy (anonymous):

yep, it's right, if you want too simplify more I think you can put it as ((sin(theta)+1)/cos(theta))^2

OpenStudy (anonymous):

Thanks all xD

OpenStudy (anonymous):

:-)

OpenStudy (anonymous):

yu have any doubt???????????

OpenStudy (anonymous):

Because I can make typos :D